A no of bullets are fired in all direction with the same intial velocity of u. the maximum area of ground coverd by bullets is

Grade:12

1 Answers

Kirti
6 years ago
Here bullets are fired in all the directions..this means that they are thrown in a circular way ...Area of circle is pie r ^ 2Here r = rangeR= u ^ 2 sin 2 theta /g R = u ^ 2 / g A = pie ×( u ^ 2 /g ) ^ 2A = pie × u ^4 / g ^ 2